A Complete Bibliography of Publications in Mathematical Tables and [Other] Aids to Computation, 1940–1949

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 January 2024 Version 2.16

Title word cross-reference

38 [Mit48]. I_{n}(X) [D. 44]. π [FW48, SWF47]. Y_{0}(z), Y_{1}(z) [HS49]. Y′_{1}(z) [HS49].

50th [Gau94].

602 [Ver49].

Accuracy [Ste49]. Addenda [Ano45b, Ano45c]. Addendum [Ano43c, Ano45d]. Admiralty [TS47]. Advanced [R. 46]. Aid [Ver49]. Aids [Ano48e, Sti48]. Airflow [Cli48, Lot48]. algebraic [Fra45]. Analog [McC49]. Analyzer [Lev48]. Anniversary [Gau94]. Application [Hae48, L. 46]. Applications [McP48]. applied [Leh45]. Approximate [Lev48]. Approximation [FW48, Har49, SWF47]. Arguments [R.C43b]. August [Gau94]. Author [Arc47]. Automatic [BCE48a, Cli48, HL48, Hof49, Hol48, KL49, Lev48, Lot48, Lub48, McP48, Mer49, Moo49, Sti47b, Ano48f, Ano48g, Ano48h, Ano48i, Ano49e, Ano49f, Ano49g, Ano49h, Lyn47, Raj47].

B.A.A.S. [R.C43a]. be [Sti47b]. Bell [Alt48a, Alt48b]. Bessel [Abr45, Ano46e, BA44, D. 44]. Bibliography [Arc47]. Binary [KL49]. Biography [Arc47]. Boundary [Hol48, Mer49, SL49]. Britain [Wom46]. British [Gau94].

Calculating [Ble47, Cur47, L. 46, Ver49]. Calculations [Mil44]. Calculator [E.48]. Calculators [Har49]. California [McC49]. Card [HC44]. Cards [W. 45]. century [Gau94]. Certain [Ano46e, Pra43, Abr45]. Class [Ano48e]. Coding [SL49]. Coefficients [Sal48b, Sal48a]. Columbia [Gau94]. Commercial [L. 46]. Comparison [Hof49]. Complex [HS49]. Compressible [Hol48]. Computation [D. 44, Hae48, Lot48, LWF72, R. 46, Gau94, Ano48e, Sti48]. computational [Gau94]. Computer [BCE48a, BCE48b, GG46, Gol72, Lyn47, McC49]. Computers [Sti47b]. Computing [Alt48a, Alt48b, Ano48e, Ano48f, Ano48g, Ano48h, Ano48i, Ano49e, Ano49f, Ano49g, Ano49h, BCE48a, Cli48, HL48, Hof49, Hol48, KL49, L. 46, Lev48, Lot48, Lub48, Lyn47, McP48, Mer49, Moo49, Raj47, Sti47b, TS47, Wom46]. concerning [Uhl46]. Conclusion [FW48]. Cone [Lot48]. Conference [R. 46]. Considerations [BCE48a]. Conversion [KL49]. Corrigenda [Ano43c, Ano43a, Ano43b, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45d, Ano45c, Ano46a, Ano46b, Ano46c, Ano46d, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d]. Cubic [Ble47]. Cumulative [LWF72]. Cylinder [Lot48].

Decimal [KL49, Lub48]. Decimals [Mil44]. Design [BCE48a, BCE48b]. Developments [Ano48f, Ano48g, Ano48h, Ano48i, Ano49e, Ano49f, Ano49g, Ano49h, Raj47]. Differential [Lev48, Mer49]. Digital [Cur47, Har49]. Discussions [BCE48a, Cli48, HL48, Hof49, Hol48, KL49, Lev48, Lot48, Lub48, McP48, Mer49, Moo49, Sti47b]. Doppler [Hof49]. Double [FM49]. Dynamics [Pra43].

EDVAC [KL49]. Edvard [Arc47]. Electric [McC49, Sti48]. Electronic [Hae48, GG46]. Elliptic [Fle48]. Engineer [Arc47]. ENIAC [GG46, Cli48]. Equation [Lev48]. Equations [Ble47, HL48, Lad48, Mer49, Moo49, Ver49, Fra45]. Equipment [Kin49]. Expressing [Sal48b, Sal48a].

Factor [Bee47]. Film [Sti47a]. Finance [Ste49]. finder [Sti48]. Finding [Moo49]. First [Sal48b, Sal48a]. Flow [Hol48]. Fluid [Hol48]. Form [KL49]. Four [Sal48b]. Fractional [Ano46e, Abr45]. Function [D. 44]. Functions [Ano46e, BA44, Bic45, Fle48, Pra43, R.C43b, Abr45, BM46].

G [Arc47]. General [BCE48a, Moo49]. Graeffe [Leh45]. Great [Wom46]. Guide [BA44, Fle48]. half [Gau94]. half-century [Gau94]. Hermite [Sal48a]. High [McP48]. High-Speed [McP48]. Hilbert [FM49].

IBM [Kin49, E.48]. II [Alt48b]. Index [LWF72]. Institute [McC49]. integral [BM46]. Integrator [GG46]. Interpolation [Ste49]. Interval [HC44]. Inversion [Che47, Mit48].

Laboratories [Alt48a, Alt48b]. Laminar [Hol48]. Laplace [SL49]. Large [Cur47, Har49, McP48, Sti47b]. Large-Scale [Har49, McP48]. Latent [FM49]. Layer [Hol48]. Legendre [Sal48b]. Linear [Lad48, Mer49, Ste49, Ver49]. Locations [Lub48]. Logical [BCE48b].

Machine [Alt48a, Alt48b, Ble47, Mer49].

Machinery [BCE48a, Cli48, Cur47, HL48, Hof49, Hol48, KL49, Lev48, Lot48, Lub48, McP48, Mer49, Moo49, Sti47b, Ano48f, Ano48g, Ano48h, Ano48i, Ano49e, Ano49f, Ano49g, Ano49h, Lyn47, Raj47]. Machines [Fra45, Hof49, L. 46, Lub48, McP48]. Mag. [R. 44]. many [Mil44]. Mathematical [R.C43a, R. 44, W. 45, EH47]. Mathematics [LWF72, Ste49, Gau94, Gau94]. Mathieu [Bic45, BM46]. Matrices [Alt46]. Matrix [FM49, Mit48]. Mechanician [Arc47]. Memory [Hae48]. Mersenne [Uhl46]. Method [Ble47, Kin49, Lad48, Moo49]. Multiplication [Alt46].

Neumann [Gol72]. Non [R.C43b].

Non-Sexagesimal [R.C43b]. Nonlinear [HL48]. Note [Che47, D. 44]. Number [Ble47, Uhl46]. Numbers [KL49]. Numerical [GG46, Mer49, Mil44].

Observations [Hof49]. Occurring [Pra43]. Odd [Ble47]. Optimum [HC44]. Optimum-Interval [HC44]. Order [Ano46e, Mit48, Abr45, BM46]. Ordinary [Mer49]. Other [Ano48e, Sti48].

P. [Arc47]. Parabolic [HL48]. Partial [Lev48]. Pascal [Gol72]. Phil. [R. 44]. Piecewise [Har49]. Places [Mil44]. Planned [Cli48]. Plotting [Kin49]. Pluggable [E.48]. Point [Lub48, Mer49]. Polynomial [Har49, Moo49]. Polynomials [Sal48b, Sal48a]. Power [Che47, Leh45]. Powers [Sal48b, Sal48a]. Principal [FM49]. Printing [EH47]. Problem [Cli48, SL49]. Procedure [Mer49]. process [Leh45]. Publicist [Arc47]. Punch [Ver49]. Punched [HC44, W. 45]. Punched-Card [HC44].

Quadratic [Ble47].

Raytheon [BCE48a, BCE48b]. Reduction [Hof49]. Relay [E.48]. Reports [R.C43a]. result [Uhl46]. Resulting [FM49]. Root [Lad48, Sti48]. Root-finder [Sti48]. Roots [FM49, Moo49]. Rule [Sti47a].

Scale [Cur47, Har49, McP48]. Scarce [Bee47]. Scheutz [Arc47]. Scientific [Arc47, L. 46, Wom46]. Sequence [E.48]. Series [Che47, FM49, Mil44, Leh45]. Service [TS47]. Sexagesimal [R.C43b]. Should [Sti47b]. Simultaneous [Lad48, Ver49]. Slide [Sti47a]. Small [Sti47b]. Solution [Ble47, Lev48, Mer49, Ver49]. Solving [Lad48, Fra45]. Speed [McP48]. Square [Lad48]. Standard [Kin49]. Statistical [McP48]. Structures [Pra43]. Symposium [Cur47, Gau94].

Table [Bee47]. Tables [BA44, Fle48, HC44, Pra43, R.C43a, R.C43b, R. 44, Ste49, W. 45, EH47]. Tabulation [Bic45, BM46]. Technical [Ano48f, Ano48g, Ano48h, Ano48i, Ano49e, Ano49f, Ano49g, Ano49h, Raj47]. Techniques [R. 46]. Technology [McC49]. Telephone [Alt48a, Alt48b]. Terms [Sal48b, Sal48a]. Testing [Mil44]. their [BM46]. Theorem [FM49]. Thirty [Sal48a]. Trigonometric [R.C43b]. Tube [Hae48]. Twenty [Sal48b]. Twenty-Four [Sal48b]. Two [Mer49]. Two-Point [Mer49].

UNIVAC [SL49]. Used [Hof49].

Value [SL49]. Values [Mer49]. Vancouver [Gau94]. Various [Hof49]. Volumes [LWF72].

Work [McP48].

Zeros [Abr45, Ano46e, HS49]. Zuse [Lyn47].

References

Abramowitz:1945:ZCB

[Abr45] Milton Abramowitz. Zeros of certain Bessel functions of fractional order. Mathematical Tables and Other Aids to Computation, 1(9):353–354, January 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Alt:1946:MM

[Alt46] Franz L. Alt. Multiplication of matrices. Mathematical Tables and Other Aids to Computation, 2(13):12–13, January 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Alt:1948:BTLa

[Alt48a] Franz L. Alt. A Bell Telephone Laboratories’ computing machine—I. Mathematical Tables and Other Aids to Computation, 3(21):1–13, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Alt:1948:BTLb

[Alt48b] Franz L. Alt. A Bell Telephone Laboratories’ computing machine—II. Mathematical Tables and Other Aids to Computation, 3(22):69–84, April 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1943:Cb

[Ano43a] Anonymous. Corrigenda. Mathematical Tables and Aids to Computation, 1(2):68, April 1943. CODEN MTTCAS. ISSN 2836-5658.

Anonymous:1943:Cc

[Ano43b] Anonymous. Corrigenda. Mathematical Tables and Aids to Computation, 1(3):100, July 1943. CODEN MTTCAS. ISSN 2836-5658.

Anonymous:1943:CA

[Ano43c] Anonymous. Corrigenda et addendum.

Mathematical Tables and Aids to Computation, 1(4):132, October 1943. CODEN MTTCAS. ISSN 2836-5658.

Anonymous:1944:Ca

[Ano44a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 1 (5):171, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1944:Cb

[Ano44b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 1 (6):204, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1944:Cc

[Ano44c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 1 (7):308, July 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1944:Cd

[Ano44d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 1 (8):336, October 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1945:C

[Ano45a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 1 (10):408, April 1945. CODEN MTTCAS.

ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1945:CAa

[Ano45b] Anonymous. Corrigenda et addenda. Mathematical Tables and Other Aids to Computation, 1(9):375, January 1945. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1945:CAc

[Ano45c] Anonymous. Corrigenda et addenda. Mathematical Tables and Other Aids to Computation, 1(12):460, October 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1945:CAb

[Ano45d] Anonymous. Corrigenda et addendum. Mathematical Tables and Other Aids to Computation, 1(11):432, July 1945. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1946:Ca

[Ano46a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2(13):63–64, January 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1946:Cb

[Ano46b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2(14):95–96, April 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1946:Cc

[Ano46c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2 (15):148, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1946:Cd

[Ano46d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2(16):196, October 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1946:MZC

[Ano46e] Anonymous. More zeros of certain Bessel functions of fractional order. Mathematical Tables and Other Aids to Computation, 2(15):118–119, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1947:Ca

[Ano47a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2(17):228, January 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1947:Cb

[Ano47b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2 (18):288, April 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1947:Cc

[Ano47c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2 (19):320, July 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1947:Cd

[Ano47d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 2(20):375, October 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:Ca

[Ano48a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3 (21):67, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:Cb

[Ano48b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3 (22):147, April 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:Cc

[Ano48c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3 (23):227, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:Cd

[Ano48d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3(24):332, October 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:NCC

[Ano48e] Anonymous. A new class of computing aids (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 3(23):217–221, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:TDAa

[Ano48f] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(21):44, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:TDAb

[Ano48g] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(22):121, April 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:TDAc

[Ano48h] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(23):206, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:TDAd

[Ano48i] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(24):317, October 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:Ca

[Ano49a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3(25):398, January 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:Cb

[Ano49b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3 (26):452, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:Cc

[Ano49c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3 (27):499, July 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:Cd

[Ano49d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 3(28):563, October 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:TDAa

[Ano49e] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(25):373, January 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:TDAb

[Ano49f] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(26):426, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:TDAc

[Ano49g] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(27):486, July 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1949:TDAd

[Ano49h] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 3(28):532, October 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Archibald:1947:PGS

[Arc47] Raymond C. Archibald. P. G. Scheutz, publicist, author, scientific mechanician, and Edvard Scheutz, engineer—biography and bibliography. Mathematical Tables and Other Aids to Computation, 2(18):238–245, April 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘R. C. A.’ as author: Goldstine [Gol72, page 16–17] confirms that the author of this paper is Raymond C. Archibald.

Bateman:1944:GTB

[BA44] Harry Bateman and Raymond Clare Archibald. A guide to tables of Bessel functions. Mathematical Tables and Other Aids to Computation, 1(7):205–308, July 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bloch:1948:GDC

[BCE48a] R. M. Bloch, R. V. D. Campbell, and M. Ellis. General design considerations for the Raytheon computer (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(24):317–323, October 1948. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bloch:1948:LDR

[BCE48b] R. M. Bloch, R. V. D. Campbell, and M. Ellis. The logical design of the Raytheon computer. Mathematical Tables and Other Aids to Computation, 3(24):286–295, October 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Beeger:1947:SFT

[Bee47] N. G. W. H. Beeger. On a scarce factor table. Mathematical Tables and Other Aids to Computation, 2(20):326–330, October 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bickley:1945:TMF

[Bic45] W. G. Bickley. The tabulation of Mathieu functions. Mathematical Tables and Other Aids to Computation, 1(11):409–419, July 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bleick:1947:CMS

[Ble47] W. E. Bleick. Calculating machine solution of quadratic and cubic equations by the odd number method. Mathematical Tables and Other Aids to Computation, 2(20):321–324, October 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bickley:1946:MFI

[BM46] W. G. Bickley and N. W. McLachlan. Mathieu functions of integral order and their tabulation. Mathematical Tables and Other Aids to Computation, 2(13):1–11, January 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Chernoff:1947:NIP

[Che47] H. Chernoff. A note on the inversion of power series. Mathematical Tables and Other Aids to Computation, 2(20):331–335, October 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clippinger:1948:APP

[Cli48] R. E. Clippinger. Airflow problem planned for the ENIAC (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(23): 206–207, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Curtiss:1947:SLS

[Cur47] J. H. Curtiss. A symposium of large scale digital calculating machinery. Mathematical Tables and Other Aids to Computation, 2(18):229–238, April 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

DHL:1944:NCB

[D. 44] D. H. L. Note on the computation of the Bessel function I_{n}(X). Mathematical Tables and Other Aids to Computation, 1(5): 133–135, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘D. H. L.’ as author: probably D. H. Lehmer.

WJE:1948:IPS

[E.48] W. J. E. The IBM pluggable sequence relay calculator. Mathematical Tables and Other Aids to Computation, 3(23):149–161, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘W. J. E.’ as author.

WJE:1947:PMT

[EH47] W. J. E. and Ralph F. Haupt. The printing of mathematical tables. Mathematical Tables and Other Aids to Computation, 2(17):197–202, January 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘W. J. E.’ as first author.

Fletcher:1948:GTE

[Fle48] Alan Fletcher. Guide to tables of elliptic functions. Mathematical Tables and Other Aids to Computation, 3(24):229–281, October 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fairthorne:1949:HDS

[FM49] R. A. Fairthorne and J. C. P. Miller.

Hilbert’s double series theorem and principal latent roots of the resulting matrix. Mathematical Tables and Other Aids to Computation, 3(26):399–400, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Frame:1945:MSA

[Fra45] J. S. Frame. Machines for solving algebraic equations. Mathematical Tables and Other Aids to Computation, 1(9):337–353, January 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ferguson:1948:NAC

[FW48] D. F. Ferguson and John W. Wrench, Jr. A new approximation to π (conclusion). Mathematical Tables and Other Aids to Computation, 3(21):18–19, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002657.

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Goldstine:1946:ENI

[GG46] H. H. Goldstine and Adele Goldstine. The Electronic Numerical Integrator and Computer (ENIAC). Mathematical Tables and Other Aids to Computation, 2(15):97–110, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goldstine:1972:CPN

[Gol72] Herman Heine Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0 (paperback), 0-691-08104-2. xii + 378 pp. LCCN TK7885.A5 G64 1993. URL http://www.

jstor.org/stable/j.ctt7rvp0. Second

printing, 1973. Paperback edition 1980.

Fifth printing, 1993 with new preface. Reprint 2000 by Books on Demand, Ann Arbor, MI, USA.

Haeff:1948:MTA

[Hae48] Andrew V. Haeff. The memory tube and its application to electronic computation. Mathematical Tables and Other Aids to Computation, 3(24):281–286, October 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Harrison:1949:PPA

[Har49] Joseph O. Harrison, Jr. Piecewise polynomial approximation for large-scale digital calculators. Mathematical Tables and Other Aids to Computation, 3(26):400–407, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Herget:1944:OIP

[HC44] Paul Herget and G. M. Clemence.

Optimum-interval punched-card tables. Mathematical Tables and Other Aids to Computation, 1(6):173–176, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hicks:1948:NPE

[HL48] Bruce L. Hicks and H. G. Landau. Nonlinear parabolic equations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(23):207–208, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hoffleit:1949:CVC

[Hof49] Dorrit Hoffleit. A comparison of various computing machines used in the reduction of Doppler observations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(25):373–377, January 1949. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Holberton:1948:LBL

[Hol48] John V. Holberton. Laminar boundary layer flow in a compressible fluid (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(23):208, July 1948. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hillman:1949:CZ

[HS49] Abraham Hillman and Iva Sherman. Complex zeros of Y_{0}(z), Y_{1}(z), and Y′_{1}(z). Mathematical Tables and Other Aids to Computation, 3(25):351–352, January 1949. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

King:1949:MPS

[Kin49] Gilbert W. King. A method of plotting on standard IBM equipment. Mathematical Tables and Other Aids to Computation, 3(25):352–355, January 1949. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Koons:1949:CND

[KL49] Florence Koons and Samuel Lubkin. Conversion of numbers from decimal to binary form in the EDVAC (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(26):427–431, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

LJC:1946:ACC

[L. 46] L. J. C. The application of commercial calculating machines to scientific computing. Mathematical Tables and Other Aids to Computation, 2(16):149–159, October 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘L. J. C.’ as author; possibly L. J. Comrie.

Laderman:1948:SRM

[Lad48] Jack Laderman. The square root method for solving simultaneous linear equations. Mathematical Tables and Other Aids to Computation, 3(21):13–16, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1945:GPA

[Leh45] D. H. Lehmer. The Graeffe process as applied to power series. Mathematical Tables and Other Aids to Computation, 1(10): 377–383, April 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Levin:1948:ASP

[Lev48] Joseph H. Levin. On the approximate solution of a partial differential equation on the differential analyzer (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(23):208–209, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lotkin:1948:CAA

[Lot48] M. Lotkin. Computation of the airflow about a cone cylinder (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(23):209–210, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lubkin:1948:DPL

[Lub48] Samuel Lubkin. Decimal point locations in computing machines (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(21):44–50, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1972:CIM

[LWF72] Yudell L. Luke, Jet Wimp, and Wyman Fair, editors. Cumulative Index to Mathematics of Computation, Volumes 1–24 (1943–1969). American Mathematical Society, Providence, RI, USA, 1972. xvii + 461 pp.

Lyndon:1947:ZCA

[Lyn47] Roger C. Lyndon. The Zuse computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 2(20):354–359, October 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002238.

McCann:1949:CIT

[McC49] G. D. McCann. The California Institute of Technology electric analog computer. Mathematical Tables and Other Aids to Computation, 3(28):501–513, October 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

McPherson:1948:ALS

[McP48] J. L. McPherson. Applications of largescale high-speed computing machines to statistical work (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(22): 121–126, April 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Meriam:1949:PMN

[Mer49] J. L. Meriam. Procedure for the machine or numerical solution of ordinary linear differential equations for two-point linear boundary values (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(28):532–539, October 1949. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Miller:1944:NTS

[Mil44] James Charles Percy Miller. Numerical testing of series by calculations to many places of decimals. Mathematical Tables and Other Aids to Computation, 1(8):309– 312, October 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Mitchell:1948:IMO

[Mit48] Herbert F. Mitchell, Jr. Inversion of a matrix of order 38. Mathematical Tables and Other Aids to Computation, 3(23):161–166, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Moore:1949:NGM

[Moo49] Edward F. Moore. A new general method for finding roots of polynomial equations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 3(27):486–488, July 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Prager:1943:TCF

[Pra43] W. Prager. Tables of certain functions occurring in dynamics of structures. Mathematical Tables and Aids to Computation, 1(4):101–103, October 1943. CODEN MTTCAS. ISSN 2836-5658.

RCA:1944:MTP

[R. 44] R. C. A. Mathematical tables in Phil.

Mag. Mathematical Tables and Other Aids to Computation, 1(5):135–141, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘R. C. A.’ as author: possibly R. C. Archibald.

RCA:1946:CAC

[R. 46] R. C. A. Conference on advanced computation techniques. Mathematical Tables and Other Aids to Computation, 2(14):65– 68, April 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘R. C. A.’ as author: possibly R. C. Archibald.

Rajchman:1947:TDA

[Raj47] Jan A. Rajchman. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 2(20):359–361, October 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

RCA:1943:MTR

[R.C43a] R.C.A. Mathematical tables in reports of the B.A.A.S. Mathematical Tables and Aids to Computation, 1(3):69–75, July 1943. CODEN MTTCAS. ISSN 2836-5658. Original journal has only ‘R. C. A.’ as author: possibly R. C. Archibald.

RCA:1943:TTF

[R.C43b] R.C.A. Tables of trigonometric functions in non-sexagesimal arguments. Mathematical Tables and Aids to Computation, 1(2):33– 44, April 1943. CODEN MTTCAS. ISSN 2836-5658. Original journal has only ‘R. C. A.’ as author: possibly R. C. Archibald.

Salzer:1948:CEFb

[Sal48a] Herbert E. Salzer. Coefficients for expressing the first thirty powers in terms of the Hermite polynomials. Mathematical Tables and Other Aids to Computation, 3(23): 167–169, July 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1948:CEFa

[Sal48b] Herbert E. Salzer. Coefficients for expressing the first twenty-four powers in terms of the Legendre polynomials. Mathematical Tables and Other Aids to Computation, 3(21):16–18, January 1948. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Snyder:1949:CLB

[SL49] Frances E. Snyder and Hubert M. Livingston. Coding of a Laplace boundary value problem for the UNIVAC. Mathematical Tables and Other Aids to Computation, 3(25):341–350, January 1949. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stelson:1949:ALI

[Ste49] Hugh E. Stelson. The accuracy of linear interpolation in tables of the mathematics of finance. Mathematical Tables and Other Aids to Computation, 3(26):408–412, April 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stibitz:1947:FSR

[Sti47a] G. R. Stibitz. Film slide rule. Mathematical Tables and Other Aids to Computation, 2(20):325, October 1947. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stibitz:1947:SAC

[Sti47b] G. R. Stibitz. Should automatic computers be large or small? (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 2(20):362–364, October 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stibitz:1948:ERF

[Sti48] G. R. Stibitz. Electric root-finder (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 3(24):328–329, October 1948. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Smith:1947:NA

[SWF47] L. B. Smith, J. W. Wrench, and D. F. Ferguson. A new approximation to π. Mathematical Tables and Other Aids to Computation, 2(18):245–248, April 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002296.

Todd:1947:ACS

[TS47] John Todd and D. H. Sadler. Admiralty Computing Service. Mathematical Tables and Other Aids to Computation, 2(19):289–297, July 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL https:/

                                              /www.ams.org/journals/mcom/1947-02-
                                              019/S0025-5718-47-99582-3;   https:
                                              //www.jstor.org/stable/2002588.

Uhler:1946:NRC

[Uhl46] Horace S. Uhler. A new result concerning a Mersenne number. Mathematical Tables and Other Aids to Computation, 2(14): 94, April 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Verzuh:1949:SSL

[Ver49] Frank M. Verzuh. The solution of simultaneous linear equations with the aid of the 602 Calculating Punch. Mathematical Tables and Other Aids to Computation, 3(27): 453–462, July 1949. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

WJE:1945:MTP

[W. 45] W. J. E. Mathematical tables on punched cards. Mathematical Tables and Other Aids to Computation, 1(12):433–436, October 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘W. J. E.’ as author.

Womersley:1946:SCG

[Wom46] J. R. Womersley. Scientific computing in great britain. Mathematical Tables and Other Aids to Computation, 2(15):110–117, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).